OpenStudy (anonymous):

Arnold went shopping at the Link with $80 to spend. He spent $12 on a belt and $20 for a shirt. How much can he spend on a pair of pants? A. x < 68 B. x < 58 C. x < 48 D. x < 38

OpenStudy (anonymous):

y = x - 32 since he already spent 32 dollars..so the answer is x should be Greater than 38 which is choice D

OpenStudy (anonymous):

Check: 38+32= 80 so that makes sence (:

OpenStudy (anonymous):

C. x < 48

OpenStudy (anonymous):

80 - 32 = 48

OpenStudy (anonymous):

oh crap lol mine equaled 70 now that i looked back :P so it's C not D

OpenStudy (anonymous):

Should you need an vision check up :P